H2O is a molecule with a bent shape. The bond angle between the two adjacent hydrogen atoms is 104.5° and the distance between them is 1.516Å. What is the length of the O-H bonds? Round your answer to 4 significant digits.

Answer by Alan3354(69443)

H2O is a molecule with a bent shape. The bond angle between the two adjacent hydrogen atoms is 104.5° and the distance between them is 1.516Å. What is the length of the O-H bonds? Round your answer to 4 significant digits.
-------------
It's an isosceles triangle.
--> a right triangle with one leg = 1.516/2 = 0.758 A and the angle opposite = 104.5/2 = 52.25 degs.
----
The O-H bond is the hypotenuse.
sin(52.25) = 0.758/hyp
hyp = 0.758/sin(52.25)
=~ 0.8278 Å